OpenStudy (anonymous):

Please help guide me through this... fan and medals! What is the solution to the system of equations represented by the two equations? y=-2x + 7 y=1/3x

OpenStudy (anonymous):

@dan815 @mathmath333 can you help?

OpenStudy (jdoe0001):

hint: \(\large { \begin{array}{llll} {\color{brown}{ y}}=-2x + 7\\ {\color{brown}{ y}}=\frac{1}{3}x \end{array}\qquad {\color{brown}{ y}}={\color{brown}{ y}}\qquad -2x + 7=\frac{1}{3}x}\) solve for "x"

OpenStudy (jdoe0001):

once you get "x", use that on either equation, to get "y:

OpenStudy (anonymous):

I'm not sure how to combine 1/3x and -2x here...

OpenStudy (anonymous):

Because wouldn't you add -2x - 1/3x +7?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well. youd' want to multiply both sides by 3 first then simplify to get "x"

OpenStudy (anonymous):

So -2x times 3 +7 = 1/3x times 3?

OpenStudy (anonymous):

Then add like terms?

OpenStudy (jdoe0001):

well \(\bf 3\cdot 2x + 3\cdot 7=\cancel{3}\cfrac{1}{\cancel{3}}x\implies 6x+21=x\)

OpenStudy (jdoe0001):

hmm actually should be -2x so

OpenStudy (anonymous):

-6x?

OpenStudy (jdoe0001):

\(\bf 3\cdot -2x + 3\cdot 7=\cancel{3}\cfrac{1}{\cancel{3}}x\implies -6x+21=x\) yeap

OpenStudy (anonymous):

so... now would you divide by -6 for both sides?

OpenStudy (jdoe0001):

nope add +6x to both sides then divide by 7

OpenStudy (anonymous):

So... 3=x?

OpenStudy (jdoe0001):

\(\bf \cancel{-6x+6x}+21=+6x+x\implies 21=7x\implies \cfrac{21}{7}=x\implies x=3\) yeap

OpenStudy (jdoe0001):

so now we know x = 3 so in " y=-2x + 7" x = 3 thus y=-2(3) + 7 y = -6+7 y = 1

OpenStudy (anonymous):

So of the other one, y=1/3(3)?

OpenStudy (anonymous):

So, (1,1) is my answer

OpenStudy (jdoe0001):

doesn't matter you could do the other if you wish BUT the other one will give the same but yes \(\bf y=\cfrac{1}{\cancel{3}}\cancel{3}\)

OpenStudy (jdoe0001):

well. recall (x,y) x = 3 and y = 1 thus -> (3,1)

OpenStudy (anonymous):

Oh, duh of course xP.

OpenStudy (anonymous):

Thank you so much!!!!